OpenStudy (anonymous):

given that f(x) =x-2 and g(x)= x + 7 find f o g

OpenStudy (anonymous):

a. x+5 b. x + 14 c. x-5 d. x-14

OpenStudy (ash2326):

\[f(x)=x-2\] and \[g(x)=x+7\] \[f o g= f(g(x))\] so \[f(g(x))=g(x)-2\] put the value of g(x) \[f(g(x))=x+7-2\] On simplification, we get \[f(g(x))=x+5\]